OpenStudy (anonymous):

solve the inequality 6x^2<6=5x

OpenStudy (danjs):

6x^2 - 5x -6 < 0

OpenStudy (danjs):

Factor ( )( )<0

OpenStudy (danjs):

(2x - 3)(3x + 2) < 0

OpenStudy (anonymous):

Thank you!

OpenStudy (danjs):

u have it from there?

OpenStudy (anonymous):

I just need to set everything equal to zero right? x= 3/2 and x=-2/3?

OpenStudy (danjs):

can you retype the original equation please?

OpenStudy (danjs):

\[x > \frac{ 3 }{ 2 } -or- x > \frac{ -2 }{ 3 } \]

OpenStudy (anonymous):

6x^2<6+5x

OpenStudy (danjs):

If you plot both those results on the number line... The entire solution can be described by just x > -2/3

OpenStudy (danjs):

You see that?

OpenStudy (anonymous):

I see it, thank you!